OpenStudy (anonymous):

Briley opens a credit card with an APR of 13.99%. How much is charged in interest this month if her balance is $2500? $349.75 $13.99 $23.32 $29.15

OpenStudy (dumbcow):

\[\frac{2500 *.1399}{12}\]

OpenStudy (anonymous):

If the APR is 13.99%, then the monthly rate is 1.1658333 % ( 13.99/12 ) A balance of $2500.00 would then costs: 2500 * (1.1658333/100 ) = $ 29.1458 or $29.15 You can also write it as: 2500 * 0.01165833 = $29.15